OpenStudy (anonymous):

a polish recipe for making 6 servings of potato salad includes 1/2 an onion. what fraction of the onion is in each serving? Show in steps please?

OpenStudy (anonymous):

I think that it is .5/6=1/12 (1/2)/6 turns into 1/2*1/6 1*1=1 2*6=12 1/12